### v2

# A polynomial bound on the number of minimal separators and potential maximal cliques in $P_{6}$-free graphs of bounded clique number

Abstract
In this note we show a polynomial bound on the number of minimal separators and potential maximal cliques in $P_{6}$-free graphs of bounded clique number.
1 Introduction
Let $G$ be a graph. For a set $X\subseteq V(G)$, we say that a connected component $C$ of $G-X$ is a full component of $X$ if $N_{G}(C)=X$, that is, every vertex of $X$ has a neighbor in $C$. A set $X\subseteq V(G)$ is a minimal separator if it has at least two full components.
Intuitively, the space of all minimal separators of a graph $G$ reflects the space of all possible separations that can be used to solve some computational problem on $G$ via dynamic programming. A related notion of potential maximal clique (not defined formally in this work) corresponds to all reasonable choices of a bag in a tree decomposition of $G$.
Bouchitté and Todinca [2] (with some results generalized by Fomin, Todinca, and Villanger [7]) showed that indeed these notions are useful to solve a wide family of graph problems.
Theorem 1 ([2, 3]).
If $G$ is an $n$-vertex graph with $a$ minimal separators and $b$ potential maximal cliques, then $b\leq n(a^{2}+a+1)$ and $a\leq nb$. Furthermore, given a graph $G$, one can in time polynomial in the input and compute the list of all its minimal separators and potential maximal cliques.
Theorem 2 (informal statement of [2, 7]).
A wide family of graph problems, including Maximum Weight Independent Set and Feedback Vertex Set, can be solved in time polynomial in the size of the input graph and the number of its potential maximal cliques.
As a result, if for a graph class $\mathcal{G}$ one can prove a (purely graph-theoretical) polynomial bound on the number of minimal separators or potential maximal cliques in $\mathcal{G}$, then one can immediately obtain polynomial-time algorithms for a wide family of problems on $\mathcal{G}$. We call such a graph class tame. Recently, a methodological study of which graph classes are tame was initiated [9, 10, 1, 14, 8].
We say that a graph $G$ is $H$-free for a graph $H$ if no induced subgraph of $G$ is isomorphic to $H$. The clique number of a graph $G$ is the maximum cardinality of a set pairwise adjacent vertices of $G$. For an integer $t$, by $P_{t}$ we denote the path on $t$ vertices (and $t-1$ edges). The main result of this note is a short proof that $P_{6}$-free graphs of bounded clique number have a polynomial number of minimal separators and potential maximal cliques.
Theorem 3.
Let $G$ be an $n$-vertex $P_{6}$-free graph of clique number $k$. Then, $G$ contains at most $(2n)^{k+1}$ minimal separators and at most $2^{2k+2}n^{2k+3}$ potential maximal cliques.
We remark that some additional condition to just $P_{6}$-freeness is needed for the conclusion of Theorem 3, as even the class of $P_{5}$-free graphs is not tame. This can be witnessed by prisms: An $n$-prism consists of two $n$-vertex cliques with a matching in between; it is $P_{5}$-free but admits $22^{n}-$ 2 minimal separators. Furthermore, the $P_{6}$ cannot be replaced with $P_{7}$: [5] provides a construction of $P_{7}$-free graphs $G_{n}$ that have clique number $2$, $|V(G_{n})|=6n+2$, and at least $3^{n}$ minimal separators.
Our motivation for proving Theorem 3 is two-fold. First, the proof is very simple, much simpler than the arguments giving polynomial-time algorithms for Maximum Weight Independent Set and related problems in $P_{6}$-free graphs without any assumption on the clique number [11]. Second, it gives examples of graph classes that are tame, but in which such problems as $5$-Coloring or Odd Cycle Transversal are NP-hard: $5$-Coloring is NP-hard in $P_{6}$-free graphs [13] (and trivial in graphs of clique number larger than $5$) and Odd Cycle Transversal is NP-hard in $P_{6}$-free graphs of clique number at most $3$ [6, 4]. This answers negatively a question of Milanič and Pivač [14, Open problem 3].
2 Proof of Theorem 3
If $G$ is edgeless, then the statement is immediate (there are no minimal separators and $n$ potential maximal cliques), so we assume $E(G)\neq\emptyset$ and thus $n,k\geq 2$.
Assuming $n,k\geq 2$, we will prove that $G$ contains at most $(2n)^{k}\cdot(2n-1)$ minimal separators. This directly implies the bound for minimal separators of Theorem 3 and also implies the bound on the number of potential maximal cliques in $G$ via Theorem 1.
We need modules and some basic facts about them. Let $G$ be a graph. A module is a nonempty set $M\subseteq V(G)$ such that every $u\in V(G)\setminus M$ is adjacent to either all vertices of $M$ or to no vertex of $M$. A module $M$ is:
  • strong if for every other module $M^{\prime}$ either $M\subseteq M^{\prime}$, $M^{\prime}\subseteq M$, or $M\cap M^{\prime}=\emptyset$;
  • proper if it is different than $V(G)$;
  • maximal if it is proper and strong and no other proper strong module contains it;
  • connected if $G[M]$ is connected.
We need the following properties of modules, which can be distilled from Lemma 2 and Theorem 2 and the discussion around them in [12].
Lemma 4.
In a graph on at least two vertices, the maximal modules form a partition of the vertex set.
Lemma 5.
An $n$-vertex graph $G$ contains at most $12n-$ 1 strong modules. Furthermore, for every module $M$ in $G$, there exists a unique strong module $M^{\prime}$ that contains $M$ and is inclusion-wise minimal with this property, and one of the following holds:
  • $M=M^{\prime}$ and $M$ is a strong module in $G$;
  • $G[M^{\prime}]$ is disconnected and $M$ is a union of at least two connected components of $G[M^{\prime}]$;
  • the complement of $G[M^{\prime}]$ is disconnected and $M$ is a union of at least two connected components of the complement of $G[M^{\prime}]$.
In the context of graphs of bounded clique number, we observe the following immediate corollary.
Corollary 6.
An $n$-vertex graph of clique number $k$ contains at most $2^{k}(2n-1)$ connected modules.
Proof.
Fix a connected module $M$ and let $M^{\prime}$ be as in Lemma 5. Then, Lemma 5 implies that $M$ is a union of some connected components of the complement of $G[M^{\prime}]$ (possibly all of them for the first option of Lemma 5). There are at most $(2n-1)$ choices for $M^{\prime}$ and at most $2^{k}$ choices which components of the complement of $G[M^{\prime}]$ form $M$, as the clique number of $G$ is $k$. ∎
We will also need the following lemma of [11].
Lemma 7 (cf. Lemma 4.2 of [11]).
Let $G$ be a graph, $X\subseteq V(G)$, and let $A$ be a full component of $X$ with $|A|>1$. Let $p,q\in A$ be any two vertices in different maximal modules of $G[A]$ (which exist by Lemma 4) that are adjacent (i.e., $pq\in E(G)$). Then, for every $x\in X$ either:
  • There exists an induced $P_{4}$ in $G$ with one endpoint in $x$ and the remaining three vertices in $A$.
  • The vertex $x$ is adjacent to $p$ or to $q$ (or to both).
  • The complement of $G[A]$ is disconnected and $N(x)\cap A$ consists of some connected components of the complement of $G[A]$.
We remark that if the complement of $G[A]$ is disconnected, then the connected components of the complement of $G[A]$ are exactly the maximal proper strong modules of $G[A]$.
We deduce the following.
Lemma 8.
Let $G$ be a $P_{6}$-free graph of clique number $k\geq 2$, let $X$ be a minimal separator in $G$, and let $A$ and $B$ be two distinct full sides of $X$. Then there exists a set $Q\subseteq A$ of size at most $k$ such that every vertex of $X\setminus N(Q)$ is complete to $B$.
Proof.
If $|A|=1$, take $Q=A$, so $X\setminus N(Q)=\emptyset$ and we are done. Assume then $|A|>1$.
If the complement of $G[A]$ is disconnected, take $Q$ to be any set consisting of one vertex from each connected component of the complement of $G[A]$. Otherwise, take $Q$ of size $2$, consisting of any two vertices of two different maximal modules of $G[A]$ that are adjacent (such two modules exist as $|A|>1$ and $G[A]$ is connected). Since the clique number of $G$ is $k$, we have $|Q|\leq\max(k,2)=k$. By Lemma 7, for every $x\in X\setminus N(Q)$ there exists a $P_{4}$ with one endpoint in $x$ and the remaining three vertices in $A$; denote it $P^{x}$. If $x$ is not complete to $B$, then, as $G[B]$ is connected and $B$ is a full component of $X$, there are $y,z\in B$ with $xy,yz\in E(G)$ but $xz\notin E(G)$. Then, $P^{x}$, prolonged with $y$ and $z$ is an induced $P_{6}$ in $G$, a contradiction. ∎
By Lemma 8, we infer that $B$ is a connected module of $G-N(Q)$. As $|Q|\leq k$, there are at most $n^{k}$ choices of $Q$. By Corollary 6, for fixed $Q$, there are at most $2^{k}(2n-1)$ choices of $B$. As $X=N_{G}(B)$, there are at most $(2n)^{k}\cdot(2n-1)$ minimal separators in $G$, as promised.
Acknowledgement.
We are grateful to Jakub Gajarský, Lars Jaffke, Paloma Thomé de Lima, Jana Masařiková, Uéverton Souza, and Michał Pilipczuk for inspiring discussions and to Martin Milanič for bringing our attention to [14, Open problem 3].

### v3

@@
 In this note we show a polynomial bound on the number of minimal separators and potential maximal cliques in $P_{6}$-free graphs of bounded clique number.
+keywords:
+potential maximal cliques, minimal separators, tame class, $P_{6}$-free graphs
 1 Introduction
@@
 A related notion of potential maximal clique (not defined formally in this work) corresponds to all reasonable choices of a bag in a tree decomposition of $G$.
-Bouchitté and Todinca [2] (with some results generalized by Fomin, Todinca, and Villanger [7]) showed that indeed these notions are useful to solve a wide family of graph problems.
-Theorem 1 ([2, 3]).
+Bouchitté and Todinca (2001) (with some results generalized by Fomin et al. (2015)) showed that indeed these notions are useful to solve a wide family of graph problems.
+Theorem 1 (Bouchitté and Todinca (2001, 2002)).
 If $G$ is an $n$-vertex graph with $a$ minimal separators and $b$ potential maximal cliques, then $b\leq n(a^{2}+a+1)$ and $a\leq nb$.
 Furthermore, given a graph $G$, one can in time polynomial in the input and compute the list of all its minimal separators and potential maximal cliques.
-Theorem 2 (informal statement of [2, 7]).
+Theorem 2 (Bouchitté and Todinca (2001); Fomin et al. (2015), informal statement).
 A wide family of graph problems, including Maximum Weight Independent Set and Feedback Vertex Set, can be solved in time polynomial in the size of the input graph and the number of its potential maximal cliques.
@@
 We call such a graph class tame.
-Recently, a methodological study of which graph classes are tame was initiated [9, 10, 1, 14, 8].
+Recently, a methodological study of which graph classes are tame was initiated, see, e.g., the work of Gartland and Lokshtanov (2023a, b); Abrishami et al.
+(2022); Milanič and Pivač (2021); Gajarský et al.
+(2022).
 We say that a graph $G$ is $H$-free for a graph $H$ if no induced subgraph of $G$ is isomorphic to $H$.
@@
 This can be witnessed by prisms: An $n$-prism consists of two $n$-vertex cliques with a matching in between; it is $P_{5}$-free but admits $22^{n}-$ 2 minimal separators.
-Furthermore, the $P_{6}$ cannot be replaced with $P_{7}$: [5] provides a construction of $P_{7}$-free graphs $G_{n}$ that have clique number $2$, $|V(G_{n})|=6n+2$, and at least $3^{n}$ minimal separators.
+Furthermore, the $P_{6}$ cannot be replaced with $P_{7}$: Chudnovsky et al.
+(2023) provide a construction of $P_{7}$-free graphs $G_{n}$ that have clique number $2$, $|V(G_{n})|=6n+2$, and at least $3^{n}$ minimal separators.
 Our motivation for proving Theorem 3 is two-fold.
-First, the proof is very simple, much simpler than the arguments giving polynomial-time algorithms for Maximum Weight Independent Set and related problems in $P_{6}$-free graphs without any assumption on the clique number [11].
-Second, it gives examples of graph classes that are tame, but in which such problems as $5$-Coloring or Odd Cycle Transversal are NP-hard: $5$-Coloring is NP-hard in $P_{6}$-free graphs [13] (and trivial in graphs of clique number larger than $5$) and Odd Cycle Transversal is NP-hard in $P_{6}$-free graphs of clique number at most $3$ [6, 4].
-This answers negatively a question of Milanič and Pivač [14, Open problem 3].
+First, the proof is very simple, much simpler than the arguments of Grzesik et al.
+(2022) giving polynomial-time algorithms for Maximum Weight Independent Set and related problems in $P_{6}$-free graphs without any assumption on the clique number.
+Second, it gives examples of graph classes that are tame, but in which such problems as $5$-Coloring or Odd Cycle Transversal are NP-hard: Huang (2016) proved that $5$-Coloring is NP-hard in $P_{6}$-free graphs (and trivial in graphs of clique number larger than $5$) and Dabrowski et al.
+(2020); Chudnovsky et al.
+(2021) proved that Odd Cycle Transversal is NP-hard in $P_{6}$-free graphs of clique number at most $3$.
+This answers negatively a question of (Milanič and Pivač, 2021, Open problem 3).
 2 Proof of Theorem 3
@@
   • connected if $G[M]$ is connected.
-We need the following properties of modules, which can be distilled from Lemma 2 and Theorem 2 and the discussion around them in [12].
+We need the following properties of modules, which can be distilled from Lemma 2 and Theorem 2 and the discussion around them in the work of Habib and Paul (2010).
 Lemma 4.
@@
 ∎
-We will also need the following lemma of [11].
-Lemma 7 (cf. Lemma 4.2 of [11]).
+We will also need the following lemma of Grzesik et al. (2022).
+Lemma 7 (cf. Lemma 4.2 of Grzesik et al. (2022)).
 Let $G$ be a graph, $X\subseteq V(G)$, and let $A$ be a full component of $X$ with $|A|>1$.
@@
 As $X=N_{G}(B)$, there are at most $(2n)^{k}\cdot(2n-1)$ minimal separators in $G$, as promised.
-Acknowledgement.
-We are grateful to Jakub Gajarský, Lars Jaffke, Paloma Thomé de Lima, Jana Masařiková, Uéverton Souza, and Michał Pilipczuk for inspiring discussions and to Martin Milanič for bringing our attention to [14, Open problem 3].
+Acknowledgements.
+We are grateful to Jakub Gajarský, Lars Jaffke, Paloma Thomé de Lima, Jana Masařiková, Uéverton Souza, and Michał Pilipczuk for inspiring discussions and to Martin Milanič for bringing our attention to Open problem 3 of  Milanič and Pivač (2021).